OpenStudy (anonymous):

5(x-6)+6=7(x+2)-6 NEED HELP PLEASE

OpenStudy (anonymous):

5(x-6)+6=7(x+2)-6 start by distributing the 5 and the 7 5*x-5*6+6=7*x+7*2-6

OpenStudy (anonymous):

distribute\[5x-30+6=7x+14-6\]combine like terms\[5x-24=7x+8\]add 24 to both sides\[5x=7x+32\]subtract 7x from both sides\[-2x=32\]divide both sides by -2\[x=-16 \]

OpenStudy (anonymous):

@Elio123 helpful? :)